OpenStudy (anonymous):

The slope of the line whose equation is 3y + 2x = 1 is

OpenStudy (anonymous):

1/3x

OpenStudy (radar):

I would put this equation in the "slope intercept" form as a start. Here is what a "slope intercept" form looks like: y=mx+b. 3y + 2x = 1 3y = -2x + 1 obtained by subtracting "2x" from both sides. Now to get just y y = (-2/3)x + 1 Got that by dividing both sides by 3. It is now proper form. m=-2/3 m is also the slope.